OpenStudy (anonymous):

Every 2 days a scientist spends 1.25 hours observing dolphins and 3.5 hours observing sea lions. At this rate, how many total hours will the scientist spend observing dolphins and sea lions in 6 days?

OpenStudy (anonymous):

@Bookworm14 @twolsey5 @jabez177

OpenStudy (twolsey5):

28.5 hours

OpenStudy (bookworm14):

so you are starting with a 2 day amount of hours 2 days = 4.75 total hours (add together the amount of hours to get 4.75) so in order to get 6 days worth you multiply 2 days by 3 = 6 days now you have to do the same thing to the other side ~4.75 * 3 = 14.25 total hours after 6 days~ now if you want to know each amount of time do the same thing accept on individual #'s ~(dolphins for 6 days) 1.25 * 3= 3.75 hours total after 6 days~ ~(sea lions for 6 days) 3.5 * 3 = 10.5 hours total after 6 days~ ~once again that is 14.25 total~ Hope this helps you out!!~ Bree :)